OpenStudy (msmr):

How do you simplify a radical? (√18ab)^3(a√3a2b +b√5ab^2)

OpenStudy (msmr):

\[(\sqrt{18ab})^{3}(a \sqrt{3a ^{2}b}+b \sqrt{5ab ^{2}})\]